OpenStudy (soahelp):

what is 8/-4 ÷ -3/9.

OpenStudy (anonymous):

-8/4 / -3/9 Flip the 2nd fraction and multiply: \(\sf -\dfrac{8}{4} \times -\dfrac{9}{3}\) Multiply the numerators and denominators together.

OpenStudy (hockeychick23):

ok so the easiest way to do this is to use *keep change change* which means we flip -3/9 to be -9/3 and multiply instead of divide. So you do -8/4*-9/3 which when simplified is -2*-3

OpenStudy (anonymous):

-8 * -9 = ? 4 * 3 = ?

OpenStudy (haseeb96):

8/-4 divide -3/9 when divide sign will be remove multiply sign will be come there and the 2nd value will be reverse it will be look like this 8/-4 multiply 9/-3